Sulfur compounds cause “off odors” in wine, so winemakers want to know the odor threshold, the lowest concentration of a compound a human nose can detect, the odor threshold for dms in trained wine testers is 25 micrograms.per liter of wine,

how sensitive are the rained noses of students? You want to estimate the mean dms odor threshold among all students and you would be satisfied to estimate the mean to within plus minus 0.1 with 90 % confidence. The standard deviation of the odor threshold for untrained noses is known to be o. = 8 micrograms per liter of wine. How large an SRS of untrained students do you need?

n = ?
Answer #1

z value at 90% = 1.645

Me = 0.1 , s = 8

ME =z *(s/sqrt(n))
0.1= 1.645 *(8/sqrt(n))
n = (1.645 * 8/0.1)^2
n = 17319